RUN 2B L1CAL MEETING MINUTES 15 November, 2005 Minutes by: Mike Mulhearn with additions by Maris Abolins Status Report Mike: No notes, was talking. I recall: - Found what appears to be a problem with the ADF configuration for real BLS signals since the major test stand changes. Philipp is now investigating. - Working on GAB L2/L3 buffer and any-instrumented-TT-over-threshold and/or term. - Have made drop in TAB/GAB library for TCC. Now any changes to "expert" tab/gab code are immediately available to TCC if appropriate. Lynn: - TAB/GAB crate fuse-panel will be side mounted. - Supply monitoring needed custom wiring due to a non-standard backplane. - Install by end-of-week, shoot for ORC before thanksgiving holiday. - L2/L3 fibers ordered, second floor patch panel installed. Jorge: - 26 sets of ATC / patch pannel/ LVDS cables - So far 2 ATC , 2 patch pannels, and 1 LVDS cable with problems. - Complete set of ATCs in hand, expect last 9 patch panels wednesday. - Anticipate to finish job by Thanksgiving. Dan: - Waiting for completion of mechanical work in rack S105 , and definition of top and bottom crate in S111 from Alan, before installing remaining ADFs. Philippe: - Preparing to support full contigent of ADFs. Todd: - New unpacker committed, examine zooms in on selected towers for comparison of precision to 2B TTs. Waiting for GAB data. Talks: Adam: See transparencies. Discussion: - No remote on/off of power supplies needed or wanted. - Do we pause the run for L1Cal2B failures during "stable ops"? To put it another way "Do we want to throw very expensive perfectly useful data on the floor?" Sabine: Please read the informative transparencies that she did not have adequate time to present. She is soliciting concerns about proposed algorithm changes. Hardware limitations mandate changes in electron algorithm. Either masks have to be abandoned or EM showers need to be indexed at right- top rather than lower-left.